Accelerator Update for April 27, 2009 - April 29, 2009 |
Monday April 27 The day shift began with Operations monitoring the Tevatron (TeV) store 7000 (115.62E30) and the Recycler stash (134.1E10), with the Antiproton Source stacking (9.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations terminated store 7000 at 1:30 PM. D0 made an access into their hall and cryo system personnel from CHL completed their cryo work following store termination. Operations put the TeV into a dry squeeze at the end of the day shift. TeV personnel checked the orbits and then Operations put the TeV into HEP shot setup at 6 PM. Operations established store 7004 at 7:20 PM with an initial luminosity of 301E30. Pbar, NuMI, MiniBooNE, and MTest took beam when available. Pbar transfers to the Recycler were completed when appropriate. Tuesday April 28 The midnight shift began with Operations monitoring store 7004 (150.4E30) and the stash (161.6E10), and with Pbar stacking (16.6E10). Operations terminated store 7004 at 4 AM. Operations established store 7005 at 5:45 AM with an initial luminosity of 258.0E30. The NuMI beam line developed a target/ BPM problem during the HEP shot setup. Controls personnel were called in to investigate. Various BPM front ends and the autotune control software were reset and/or rebooted. The problem was resolved prior to shutting the beam off for the scheduled MI access. Shutdown activities began at 7 AM with accesses to Linac, Booster, MI, MI-8, pre-vault, and the vault. Operations established beam to Linac by early afternoon. Problems with the 400 MeV line Debuncher held off the Booster beam. Pbar held on to a stack throughout the day shift. The MI access finished up shortly after the evening shift began. Checkout was underway by 5:15 PM. Operations established a low intensity beam to Pbar, NuMI, and MiniBooNE by 6:15 PM. Beam was delivered to MTest from 6:40 - 10:40 PM. Proton Source personnel conducted machine tuning until late in the evening shift. Pbar transfers to the Recycler were completed when appropriate. Wednesday April 29 The midnight shift began with TeV cooldown in progress, with Operations monitoring the stash, and with Pbar stacking. Operations terminated store 7005 at 1:20 AM. Operations put the TeV into HEP shot setup and established store 7006 at 3:15 AM with an initial luminosity of 197.19E30. Pbar, NuMI, and MiniBooNE took beam when available. Pbar transfers to the Recycler were completed when appropriate. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/29/2009 email Fermilab |
| Security, Privacy, Legal | |